Mrs. Crews bought 4 pencils and 3 pens for $5.60... Mrs Henderson bought 2 pencils and 3 pens for $4.60 of the same kind... What is the price for each pen and pencil?
Answer by mananth(16946)

x= price of pencil
y= price of pen

4.00 x + 3.00 y = 5.60 .............1

2.00 x + 3.00 y = 4.60 .............2
Eliminate y
multiply (1)by -1.00
Multiply (2) by 1.00
-4.00 x -3.00 y = -5.60
2.00 x + 3.00 y = 4.60
Add the two equations
-2.00 x = -1.00
/ -2.00
x = .50
plug value of x in (1)
4.00 x + 3.00 y = 5.60
2.00 + 3.00 y = 5.60
3.00 y = 5.60 -2.00
3.00 y = 3.60
y = 1.20
x= 0.50 price of pencil
y= 1.20 price of pen
m.ananth@hotmail.ca